OpenStudy (ashking1):

Use the arc length formula and the given information to find s. r = 20 ft θ = 19° s = ?

OpenStudy (anonymous):



OpenStudy (anonymous):

first you need to convert \(19^\circ\) to the usual radian measure do you know how to do that?

OpenStudy (anonymous):

Use the arc length formula, \(S=\theta \times \frac{ \pi }{ 180 }\times r\), where \(r\) is the radius. So with the given, find S. \(\huge S=\ 19\times \frac{ \pi }{ 180 }\times 20\)

OpenStudy (jack1):

agree ;)